# Deep learning in a bilateral brain with hemispheric specialization

### 05 Mar 2024 v7

@@ Abstract
The brains of all bilaterally symmetric animals on Earth are divided into left and right hemispheres. The anatomy and functionality of the hemispheres have a large degree of overlap, but there are asymmetries and they specialize to possess different attributes. Several computation models mimic hemispheric asymmetries with a focus on reproducing human data on semantic and visual processing tasks. In this study, we aimed to understand how dual hemispheres in a bilateral architecture interact to perform well in a given task. We propose a bilateral artificial neural network that imitates lateralization observed in nature: that the left hemisphere specializes in specificity and the right in generalities. We used different training objectives to achieve the desired specialization and tested it on an image classification task with two different CNN backbones – ResNet and VGG. Our analysis found that the hemispheres represent complementary features that are exploited by a network head which implements a type of weighted attention. The bilateral architecture outperformed a range of baselines of similar representational capacity that don’t exploit differential specialization, with the exception of a conventional ensemble of unilateral networks trained on a dual training objective for specifics and generalities. The results demonstrate the efficacy of bilateralism, contribute to the discussion of bilateralism in biological brains and the principle may serves as an inductive bias for new AI systems.
Impact statement A striking feature of the brain is that it is divided into two hemispheres. It is a conserved feature across all bilaterally symmetric animals, strongly suggesting its importance for intelligence. Many researchers have investigated observed hemispheric asymmetry. We take a different approach, assuming specialization, which enables an investigation into how dual hemispheres complement each other to improve performance. The study constitutes a novel approach to understanding the role of dual hemispheres for the cognitive sciences and the analysis suggests possible mechanisms. In addition, the architecture outperformed a variety of comparable baselines, demonstrating that bilateralism is an important biologically-inspired principle to be considered in the design of new AI systems.
Keywords Hemispheric specialization, Hemispheric asymmetry, Brain-inspired architecture, Bilateral, Computational cognitive neuroscience, Deep learning, Inductive bias
@@ 1 Introduction
Many advances in the field of AI have been inspired by the human brain including the perceptron, reinforcement learning and imitation learning. In this study, we took inspiration from the bilateral nature of the brain and applied it to a standard task suited to a bilateral architecture.
What is bilateralism? The brain is divided into left and right hemispheres. It’s a remarkably conserved feature across species, suggesting its importance for intelligence. The anatomy and functionality of the hemispheres have a large degree of overlap, but they are asymmetric and specialize to possess different attributes.
A likely explanation for the emergent specialization is hemispheric asymmetry through small differences in functional and anatomical properties of the cortex. For example, neuron density distribution. The left resembles a pyramid with a greater density of neurons near input layers, whereas the right resembles an inverted pyramid, with greater neuron density near output layers [1, 2, 3, 4, 5, 6]. Network topology also plays a role. The left has more short local pathways constituting a modular network organization, whereas the right has more longer inter-regional pathways [4, 5, 6]. More differences are described in Section 2.1.1, in the Related Work.
Many researchers have used computational models to study the bilateral brain. They mimic anatomical and functional asymmetries, subject the system to standard tasks, and compare the results to human behavioral data. The various studies provide evidence that asymmetric functional and anatomical differences can give rise to observed lateralization of activity (i.e. specialization) for specific tasks. More detail is given in Section 2.1.2, in the Related Work.
The central question of the existing models, is to understand what gives rise to observed specialization. The aim of this project is to investigate a different question – what is the benefit of having differentially specialized hemispheres and how do they complement each other for a given task? Our approach is to induce specialization directly, rather than test if it emerges as a result of architectural asymmetry. We investigated bilateralism through a computational model, with the objective of developing a theory of operation and to understand how bilateral principles could be exploited for AI/ML.
@@ 1.1 Our approach
As a starting point, we focused on the left hemisphere’s specialization on specifics, and the right’s specialization on generalities [4, 5, 6]. We hypothesize that an ANN with differentially spec ialized sub-networks would outperform a single comparable network in a classification task featuring both fine and coarse class labels.
We chose a hierarchical image classification task where each sample belongs to both a fine and a coarse class. For example, coarse: sea creature, fine: penguin, seal, shark etc. We modelled the hemispheres with left and right CNNs and induced specialization using supervised training with different training objectives, using either coarse or fine labels. The bilateral network was compared to several comparative baselines that do not have specialization but do have the same number of learnable parameters. The differences were analyzed to explore the effects of bilateral specialization. Experiments were repeated across two types of CNN hemispheres, VGG and ResNet variants, to test generalization over different core CNN architecture.
We predicted that with specialization, the hemispheres would extract distinct but overlapping features, despite identical observations. In addition, we expected that a bilateral architecture would be able to combine these features to improve overall accuracy.
@@ 2 Related work
2.1 Biological computational models
Multiple studies examine hemispheric function in standard behavioral tasks. The studies use bilateral artificial neural networks that mimic functional and architectural hemispheric asymmetries. They replicate aspects of human behavioral data on the tasks, focussing on lateralized hemispheric activity. Some of the studies also consider neurological development, disorders or damage and recovery. With the exception of face perception studies, they all used synthetic abstracted representations for input and output vectors and do not operate on realistic sensory data.
This section is divided into two subsections: a) ‘Hemispheric asymmetries’ which provides additional background on observed architectural and anatomical asymmetries, and b) ‘Models and tasks’ which describes the models that were created, organized by the tasks that were used for testing them.
2.1.1 Hemispheric asymmetries
In addition to the examples given in the introduction, topology and neuron density distribution, other architectural differences come into play. From a connectivity perspective, there are different intra-hemispheric patterns of connectivity between regions within each hemisphere [8, 9]. Furthermore, inter-hemispheric communication by the Corpus collosum results in a competitive process between the hemispheres [2, 1, 10, 11] which is believed to support lateralization of activity. Asymmetry in the visual stream is caused by a split-fovea, where each hemisphere receives the contra-lateral half of the visual field [12, 13, 14, 11, 15, 16, 17].
There are also functional asymmetries. Neurotransmitters such as dopamine and norepinephrine are asymmetrically distributed and neuron firing thresholds differ between hemispheres [10]. Also, the hemispheres respond asymmetrically to frequency: the left responds to higher frequencies than the right [14, 15, 12]. Finally, there may be additional asymmetries that are not yet discovered.
2.1.2 Models and tasks
The behavioral tasks can be divided into two major groups, semantic processing and visual perception. For semantic processing, behavioral data and patterns of hemispheric asymmetry were reproduced by varying intra-hemispheric connectivity [9] in the context of semantic dementia [18], and in the context of unilateral and bilateral lesions [8]. Other studies varied anatomical features such as inter-hemispheric connectivity and the distribution of units between layers, as well as functional features such as learning rate, sensitivity and maximum unit activation and combinations of these factors [2, 1, 10] and in the context of lesions and recovery [3]. Another anatomical feature that was explored is the split-fovea [12] and in the context of dyslexia [13]. In the case of the split-fovea experiments, hemispheric asymmetry emerged as a result of the orthographic to semantic patterns inherent in the language.
Visual perception tasks can be further divided into face perception and the line-bisection task. Face perception includes identification of faces and emotions. In the line-bisection tasks, subjects are asked to mark the center of a horizontal line, as perception of the center is affected by hemispheric function. Face perception was first studied by Dailey et al. Dailey and Cottrell [15] with a model that mimicked the hemispheric frequency response asymmetry. Hsiao et al. Hsiao et al. [14] added a split-fovea and then Wang et al. Wang and Cottrell [11] incorporated other functional parameter differences between the hemispheres as well as a learned gating mechanism for inter-hemispheric communication. For the line bisection task, Shillcock et al. Shillcock and Monaghan [16] and Monaghan and Shillcock Monaghan and Shillcock [17] used a split-fovea, asymmetric frequency response as well as asymmetric distribution of units in layers between left and right.
In the closest work to this project, Mayan et al. Mayan et al. [19] tackled the same problem of image classification with fine and coarse (hierarchical) classes. The authors created a single network with two hemispheres, and trained it with supervised learning and a single loss function. They used hyperparameters with analogies to biological parameters in each hemisphere to encourage specialization. Specialization was achieved in each side but the bilateral network did not have an advantage over baselines.
2.2 Biologically inspired multi-network architectures
@@
Several studies cover related ideas, although they did not explicitly model bilateral hemispheres. Beaulieu et al. [20] created a dual-stream architecture called Neuromodulator Meta-learner, where one network learns to modulate the other, to enhance continual learning. Bakhtiari et al. [21] created a network with parallel pathways to reproduce the functionality of dorsal (‘where’) and ventral (‘what’) pathways in an ANN trained with a single loss function. Li and Deza [22] discussed specialization of a branched neural network when trained on a Gabor filter dataset and found that the training curriculum is inconsequential to specialization of the branched neural network.
2.3 Ensemble models
@@
The dual-hemisphere architecture in this work can be viewed as a type of ensemble model. Ensembling is a popular approach in ML [23]. The most common techniques involve combining the outputs of the same model type, trained with different seeds. However, there are also many techniques that encourage diversity within the ensemble using different training objectives, sampling, architecture, or losses [24, 25, 26]. Our work may be seen as a special case of a diversified ensemble, where the ensemble architecture exploits hierarchical data labels to achieve the diversification.
2.4 Multi-domain learning
@@
Another related area is multi-domain learning. In multi-domain learning (MDL), multiple models are trained to specialize on a particular domain [27]. Two of the main approaches include a) modelling domain-specific and domain-general features [28] and b) explicitly modelling the relationship between domains. MDL is focussed on classes from domains with different distributions. The domain is sometimes explicit, and other times it is determined by the model [29]. In contrast, this project is focussed on explicit hierarchical labels from the same distribution. The objective being to mimic learning different aspects of a single task with left and right hemispheres (2 separate models).
2.5 Hierarchical image datasets
We focused on image classification of fine and coarse classes, exploiting the hierarchical nature of CIFAR-100 [30]. CIFAR-100 and other hierarchical image classification datasets are widely used as benchmarks, including ImageNet [31] and Omniglot [32]. However, the hierarchical nature of available classes is rarely exploited to enhance training or generate a more difficult task. Notable exceptions arise in the Continual Few-Shot Learning literature, such as the Meta-Dataset of Triantafillou et al. [33] which generates class-hierarchy-aware episodes of data and the Tiered-ImageNet of Caccia et al. [34] which uses class hierarchy to generate Out of Distribution (OoD) data.
@@ 3 Models
We implemented a bilateral model as well as several baselines. Each model consists of one or more feature extractors, representing hemispheres, as well as one or two classifier heads. The hemispheres were implemented with two different CNN architectures to test if the effects generalize across architecture types: ResNet [35] and VGG [36]. We chose VGG and ResNet because they are common, well understood and amongst the best performing on many vision tasks.
In the case of the ResNet-based models, we took the concept of residual layers and skip connections, and empirically optimized the number of layers on the dataset, resulting in a 9-layer network shown in Figure 2, and hereafter referred to as ResNet-9.
For the VGG based model, we chose the VGG-11A variant [36], hereafter referred to as VGG-11, as it is closest in depth to ResNet-9 and has been shown to work well for CIFAR100 [37]. We used the pre-defined PyTorch implementation, but removed the final fully-connected layer as we later add our own classifier (explained below). In shorthand, the full architecture is: conv64-pool-conv128-pool-conv256-conv256-pool-conv512-conv512-pool-conv512-conv512-pool-fc4096-fc4096. ConvX = convolutional with X filters, pool = max pool, fcX = fully connected with X units. All filters are 3x3 pixels. Although VGG-11 and ResNet-9 are both convolutional networks with a similar depth, there are substantial architectural differences. VGG does not have skip connections, and there are a lot more parameters due to the fully-connected layers.
@@
All networks were trained with a cross-entropy loss function, which includes a softmax operation (not shown explicitly in the diagrams as it is part of the loss function). Details about training and evaluation are given in Section 4.1.2.
The models are depicted in Figure 1 and described in the subsections below.
Figure 1: The range of models. Feature extractors are either ResNet-9 or VGG-11. The classifier heads are single layer fully connected networks, with output dimensions the same size as the number of classes. For brevity, if not specified, the network is unspecialized.
3.1 Bilateral network - specialized
Training and assembling the bilateral network consisted of two phases. In Phase 1, individual hemispheres were trained to be specialized feature extractors with a single classifier head, Figure 1a, referred to as ‘Unilateral - specialized’. The left hemisphere was specialized for specific classes through supervised training using the fine labels. The right hemisphere was specialized for general classes through supervised training using the coarse labels. See Section 4.1.2 for details on training and evaluation.
Then in Phase 2, the hemispheres were brought together into a broader architecture consisting of two hemispheres, which serve as feature extractors, and two newly initialized classifier heads, one each for fine and coarse labels, Figure 1b, referred to as ‘Bilateral - specialized’. The weights of the individual hemispheres were frozen and the heads were trained to use the features for classifying fine and coarse labels. A single loss was calculated by adding the loss from each head.
Figure 2: ResNet-9 architecture. A convolutional neural net with 9 layers and skip connections. The format of the description within each convolutional layer is filter w x h, number of filters, stride and padding. BN = Batch Norm, MP = Max Pool and AP = Average Pool.
@@ 3.2 Baseline models
Unilateral network with specialization
To understand the effect of specialization on performance, we compared to the specialized single-hemispheres, Figure 1a, as the first baseline. Specialization was achieved using supervised training with selective labels as described above in Section 3.1, ‘Bilateral network - specialized’.
Unilateral network without specialization
To better understand the single-hemisphere, and the role of using complementary loss functions in the absence of dual hemispheres, we also compared to a single-hemisphere without specialization, Figure 1d. The single hemisphere was trained on both types of labels simultaneously, by using two classifier heads and adding the loss from both.
Bilateral network without specialization
To better understand the role of specialization, we compared the bilateral network to an equivalent network without specialization, Figure 1c. This baseline has the same number of trainable parameters. It can be viewed as a type of ensemble, where the models are combined in the same way as the specialized bilateral network.
We trained the entire network (two hemispheres and two heads) without first explicitly inducing specialization in the individual hemispheres. Like for the unilateral network and bilateral network with specialization, the loss from each head are added together to create a single loss value.
@@
The bilateral network is a type of ensemble. Ensemble learning denotes combining different models or algorithms to obtain improved performance over a single model [23]. To understand the differences between differential specialization and conventional ensembling, we compared to a 2-model ensemble, Figure 1e. We chose 2 models, so that the ensemble has approximately the same number of trainable parameters as the bilateral network. We used two unilateral models without specialization, trained from different seeds. The output of the ensemble (in training and inference) was the average output from the heads: for the fine classification, it was the average of the fine classifier heads, and for the coarse classification, it was the average of the coarse classifier heads. In order to test multiple seeds, we trained five models with different seeds, and then constructed five 2-model ensembles, each with a unique pair of models selected from the pool of five.
@@ 4.1.1 Dataset
We used the CIFAR-100 dataset [30], as it includes hierarchical labels that denote fine (specific) and coarse (general) class labels. The dataset is split into 50,000 training and 10,000 test samples. Images were resized to 32x32 and we used two widely adopted data augmentation techniques, a random crop (using the RandomCrop PyTorch function) and a random horizontal flip (using the RandomHorizontalFlip PyTorch function) to increase the diversity of the training set.
@@ 4.1.2 Training and evaluation
The models were trained with supervised learning using a cross-entropy loss function. Depending on the model, the fine or coarse labels were used. In the case of dual heads (one with each label type, see Section 3), the loss from each head was added to create a single loss value. Weights were initialized with PyTorch defaults, from a random uniform distribution $U(-\sqrt{k},\sqrt{k})$: for linear layers $k=1/in\_features$, and for the convolutional layers $k=\frac{1}{C_{in}*\Pi_{i=0}^{1}kernel\_size[i]}$. We used weight decay of 1.0e-5, dropout of 0.6 on the classifier heads [38] to regularize the models, Adam [39] for optimization, with a learning rate of 1.0e-4 and a mini-batch size of 256.

### 15 Mar 2024 v8

@@ Abstract
The brains of all bilaterally symmetric animals on Earth are divided into left and right hemispheres. The anatomy and functionality of the hemispheres have a large degree of overlap, but there are asymmetries and they specialise to possess different attributes. Other authors have used computational models to mimic hemispheric asymmetries with a focus on reproducing human data on semantic and visual processing tasks. We took a different approach and aimed to understand how dual hemispheres in a bilateral architecture interact to perform well in a given task. We propose a bilateral artificial neural network that imitates lateralisation observed in nature: that the left hemisphere specialises in specificity and the right in generality. We used different training objectives to achieve the desired specialisation and tested it on an image classification task with two different CNN backbones – ResNet and VGG. Our analysis found that the hemispheres represent complementary features that are exploited by a network head which implements a type of weighted attention. The bilateral architecture outperformed a range of baselines of similar representational capacity that don’t exploit differential specialisation, with the exception of a conventional ensemble of unilateral networks trained on a dual training objective for specifics and generalities. The results demonstrate the efficacy of bilateralism, contribute to the discussion of bilateralism in biological brains and the principle may serves as an inductive bias for new AI systems.
Keywords Bilateral  $\cdot$ Hemispheric specialisation  $\cdot$ Hemispheric asymmetry  $\cdot$ Brain-inspired architecture  $\cdot$ Cognitive architectures  $\cdot$ Deep learning  $\cdot$ Computational cognitive neuroscience  $\cdot$ Inductive bias
@@ 1 Introduction
Left and right hemispheres are a remarkably conserved feature of brains across species, suggesting their importance for intelligence. The anatomy and functionality of the hemispheres have a large degree of overlap, but they are asymmetric and specialise to possess different attributes. For example, the left hemisphere is specialised for specific (fine) classes and familiarity/routine whereas the right hemisphere is specialised for more general (coarse) classes and novelty [4, 5].
Researchers have studied the bilateral brain with computational models that mimic anatomical and functional asymmetries. The models are typically subjected to standard tasks and compared to human behavioural data. The various studies provide evidence that asymmetric functional and anatomical differences can give rise to observed lateralisation of activity (i.e. specialisation) for specific tasks e.g. [1, 18, 13, 8, 11].
The central question of existing models is to understand the cause of the observed specialisation. The aim of this project is to investigate a different question – what is the benefit of specialised hemispheres and how do they complement each other for a given task? Our approach is to induce specialisation directly, rather than test if it emerges as a result of architectural asymmetry. Our objective is to develop a theory of operation and to understand how bilateral principles could be exploited for AI/ML.
@@ 1.1 Our approach
We focused on the left hemisphere’s specialisation on specifics, and the right’s specialisation on generalities [4, 5]. We hypothesised that an ANN with differentially specialised sub-networks would outperform a single comparable network in a classification task featuring both fine and coarse class labels. We predicted that with specialisation, the hemispheres would extract distinct but overlapping features, despite identical observations. In addition, we expected that a bilateral architecture would be able to combine these features to improve overall accuracy.
We chose a hierarchical image classification task where each sample belongs to both a fine and a coarse class e.g. coarse: sea creature, fine: penguin, seal, shark. We modelled the hemispheres with left and right CNNs and induced specialisation using supervised training with different objectives, using either coarse or fine labels. The bilateral network was compared to several comparative baselines that do not have specialisation but do have the same number of learnable parameters. The differences were analysed to explore the effects of bilateral specialisation. Experiments were repeated with two types of CNN hemispheres, VGG and ResNet variants, to test generalisation over different backbones.
@@ 2 Related work
Biologically inspired multi-network architectures
@@
Beaulieu et al. [20] created a dual-stream architecture called Neuromodulator Meta-learner, where one network learns to modulate the other, to enhance continual learning. Bakhtiari et al. [21] created a network with parallel pathways to reproduce the functionality of dorsal (‘where’) and ventral (‘what’) pathways in an ANN trained with a single loss function. Li and Deza [22] discussed specialisation of a branched neural network.
In the most related study, Mayan et al. [19] tackled the same image classification task with fine and coarse (hierarchical) classes. The authors created a single network with two hemispheres, and trained it with supervised learning and a single loss function. They used hyperparameters with analogies to biological parameters in each hemisphere to encourage specialisation. Specialisation was achieved in each side but the bilateral network did not have an advantage over baselines.
Ensemble models
@@
Our bilateral approach can be viewed as a type of ensemble, which is a popular approach in ML [23]. Typically, the outputs of multiple models, trained on different seeds, are combined; and the models are of the same type. Other techniques encourage diversity within the ensemble using different training objectives, sampling, architecture, or losses [24, 25, 26]. Our work can be viewed as a special case of a diversified ensemble, where the ensemble architecture exploits hierarchical data labels to achieve diversification.
Multi-domain learning
@@
In multi-domain learning (MDL), multiple models are trained to specialise in a particular domain [27]. Two of the main approaches include a) modelling domain-specific and domain-general features [28] and b) explicitly modelling the relationship between domains. MDL is focused on classes from domains with different distributions. The domain is sometimes explicit and other times it is determined by the model [29]. In contrast, this project is focused on explicit hierarchical labels from the same distribution. The objective being to mimic learning different aspects of a single task with left and right hemispheres (2 separate models).
@@ 3 Models
We compared a bilateral model to several baselines. Each model consisted of one or more feature extractor, representing hemispheres, as well as one or two classifier heads, one for each label type (fine or coarse). The hemispheres were implemented with two different CNN architectures to test if the effects generalise across architectures: ResNet [35] and VGG [36]. We chose VGG and ResNet because they are common, well understood and amongst the best performing on many vision tasks.
In the case of ResNet-based models, we empirically optimised the number of layers on the dataset, resulting in a 9-layer network, hereafter referred to as ResNet-9. In shorthand, the full architecture is: conv64-conv128-pool-res128,128-pool-conv256-conv512-pool-res512,512-av_pool. ConvX = convolutional with X filters, res = residual block with two convolutional blocks of X,Y filters, pool = max pool, av_pool = average pool. All filters are 3x3 and we used ReLU activation function.
For the VGG model, we chose the VGG-11A variant [36], hereafter referred to as VGG-11, as it is closest in depth to ResNet-9 and works well for CIFAR-100 [37]. We used the PyTorch implementation, but removed the final fully-connected layer as we added our own classifier (explained below). In shorthand, the full architecture is: conv64-pool-conv128-pool-conv256-conv256-pool-conv512-conv512-pool-conv512-conv512-pool-fc4096-fc4096. Definitions as above, and fcX = fully connected with X units. All filters are 3x3. Although VGG-11 and ResNet-9 are both convolutional networks with a similar depth, there are substantial architectural differences. VGG does not have skip connections, and there are a lot more parameters due to the fully-connected layers.
@@
The models are depicted in Figure 1 and described below. Details on training and evaluation in Section 4.1.2.
Figure 1: The range of models. Feature extractors are either ResNet-9 or VGG-11. The classifier heads are single layer fully connected networks, with output dimensions the same size as the number of classes. For brevity, if not specified, the network is unspecialised.
3.1 Bilateral network - specialised
Assembling and training the network consisted of two phases. In Phase 1, individual hemispheres were trained to be specialised feature extractors using supervised learning with a single classifier head, resulting in ‘Unilateral - specialised’, Figure 1a. The left was specialised for specific classes using fine labels, and the right for general classes with coarse labels. In Phase 2, the hemispheres (serving as feature extractors) were brought together into a broader architecture, with two initialised classifier heads, one for fine and coarse labels, resulting in ‘Bilateral - specialised’, Figure 1b, . The weights of the individual hemispheres were frozen and the heads were trained to use the features for classifying fine and coarse labels. A single loss was calculated by adding the loss from each head.
@@ 3.2 Baseline models
Unilateral networks with specialisation
We compared the bilateral network to the individual specialised hemispheres alone, on the tasks for which they were specialised, referred to as Unilateral-specialised, Figure 1a. Specialisation was achieved using supervised training with selective labels as described in Section 3.1.
Unilateral network without specialisation
To investigate specialisation, we compared Unilateral-specialised to a single hemisphere without specialisation, referred to as Unilateral-unspecialised, Figure 1d. The single hemisphere was trained on both types of labels simultaneously, by using two classifier heads and adding the loss. This baseline also helps to understand the effect of complementary loss functions; on a single hemisphere compared to the bilateral network.
Bilateral network without specialisation
To better understand the role of specialisation in the bilateral network, we compared Bilateral-specialised to an equivalent network, with the same number of trainable parameters, but without specialisation, referred to as Bilateral-unspecialised, Figure 1c. It can be viewed as a type of ensemble, where the models are combined in the same way as the specialised bilateral network.
We trained the entire network (two hemispheres and two heads) without first explicitly inducing specialisation in the individual hemispheres. Like for the unilateral network and bilateral network with specialisation, the loss from each head are added together to create a single loss.
@@
To understand the differences between differential specialisation and conventional ensembling, we compared to a 2-model ensemble, Figure 1e. We chose 2 models, so that the ensemble has approximately the same number of trainable parameters as the bilateral network. We used two unilateral models without specialisation, trained from different seeds. The output of the ensemble was the average from the heads: for fine classification, the average of the fine heads, and for coarse, the average of the coarse heads. We trained five models with different seeds, and then constructed five 2-model ensembles, each with a unique pair of models selected from the pool of five.
@@ 4.1.1 Dataset
We used the CIFAR-100 dataset [30], as it includes hierarchical labels that denote fine (specific) and coarse (general) classes. The dataset is split into 50,000 training and 10,000 test samples. Images were resized to 32x32 and we used two widely adopted data augmentation techniques, a random crop and a random horizontal flip (using the inbuilt PyTorch functions) to increase the diversity of the training set.
@@ 4.1.2 Training and evaluation
The models were trained with supervised learning using a cross-entropy loss function. In the case of dual heads (one with each label type, see Section 3), the loss from each head was added to create a single loss value. Weights were initialised with PyTorch defaults, from a random uniform distribution $U(-\sqrt{k},\sqrt{k})$: for linear layers $k=1/in\_features$, and for the convolutional layers $k=\frac{1}{C_{in}*\Pi_{i=0}^{1}kernel\_size[i]}$. We used weight decay of 1.0e-5, dropout of 0.6 on the classifier heads to regularise the models, Adam [39] for optimisation, with a learning rate of 1.0e-4 and a mini-batch size of 256.